Two Die, are rolled, what is the probability that the total number will be double numbers?
1
Expert's answer
2011-06-07T05:31:20-0400
Let’s consider the probability of just getting two k’s.
P(k, k) = P(k) · P(k) = (1/6)2. It is the same chances for all number doubles.
Therefore, the total probability for any double is:
6* (1/6)2=1/6